OpenStudy (anonymous):

The area of a rectangle is 120 cm2. The length of the rectangle is eight less than the twice of the width. Find the dimensions of the rectangle.

OpenStudy (anonymous):

so far i got....\[(2x - 8) *w=120 \right?\]

OpenStudy (anonymous):

x=w

OpenStudy (anonymous):

let width=x length=2x-8 x(2x-8)=120

OpenStudy (anonymous):

ok so i was right

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so... \[2x ^{2} - 8x -120=0\]

OpenStudy (anonymous):

carry on

OpenStudy (anonymous):

hmnmmmm

OpenStudy (anonymous):

divide by 2

OpenStudy (anonymous):

solve further

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what can i add to get 4 that i can multiply to get 60? hmm

OpenStudy (anonymous):

10 and 6

OpenStudy (anonymous):

x-6, and x-10?

OpenStudy (anonymous):

1*- 60=-60 10-6=4 10*-6=-60

OpenStudy (anonymous):

x-6, and x-10? ?

OpenStudy (anonymous):

\[x^2-\left( 10-6 \right)x-60=0\] \[x^2-10x+6x-60=0\]

OpenStudy (anonymous):

thats what i got

OpenStudy (anonymous):

\[x(x-10) 6(x-10)\]

OpenStudy (anonymous):

good

OpenStudy (anonymous):

6 cm and 20 cm 10 cm and 12 cm 40 cm and 30 cm 5 cm and 24 cm

OpenStudy (anonymous):

oh multiply by 2

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

12 and 20?

OpenStudy (anonymous):

x(x-10)+6(x-10)=0 (x-10)(x+6)=0 x+6=0,x=-6 rejected x-10=0,x=10 width=10 length=2*10-8=20-8=12

OpenStudy (anonymous):

where does 8 come from?

OpenStudy (anonymous):

oh i see

OpenStudy (anonymous):

kinda?

OpenStudy (anonymous):

in the statement length is 8 less than twice width